# Impacts of dynamic radiation transport to one-cell inversions on modern GPUs with higher-order discretization

### v1

@@ Abstract
To find deterministic solutions to the transient discrete-ordinates neutron-transport equation, source iterations (SI) are typically used to lag the scattering (and fission) source terms from subsequent iterations. For Cartesian geometries in one dimension, SI is parallel over the number of angles but not spatial cells; this is a disadvantage for many-core compute architectures like graphics processing units. One-cell inversion (OCI) is a class of alternative iterative methods that allow space-parallel radiation transport on heterogeneous compute architectures. For OCI, previous studies have shown that, in steady-state computations, spectral radius tends to unity when cells are optically thin regardless of the scattering ratio. In this work, we analyze how the convergence rate of an OCI scheme behaves when used for time-dependent neutron transport computations. We derive a second-order space-time discretization method from the simple corner balance and multiple balance time discretization schemes and show via Fourier analysis that it is unconditionally stable through time. Then, we derive and numerically solve the Fourier systems for both OCI and SI splittings of our discretization, showing that small mean-free times improve the spectral radius of OCI more than SI, and that spectral radius for OCI tends to zero as mean free time gets smaller. We extend both solvers to be energy dependent (using the multigroup assumption) and implement on an AMD MI250X using vendor-supplied batched LAPACK solvers. Smaller time steps improve the relative performance of OCI over SI, and, even when OCI requires more iterations to converge a problem, those iterations can be done much faster on a GPU. This leads to OCI performing better overall than SI on GPUs.
@@ 1 Introduction
Simulating transient particle transport is often required when computing the solution to a number of multi-physics problems, including burst criticality experiments, fission reactor accidents, and other highly dynamic systems. Finding deterministic solutions to the transient neutron transport equation requires some method of treating the contribution of scattering described by an integral. This is either done by taking moments of the neutron transport equation and making a closure assumption, or by using quadrature to discretize the integral over angle. The latter is called the method of discrete ordinates (or SN method, where $N$ is the number of angles in the quadrature) that forms a coupled set of simultaneous PDEs, with one for every direction in a given quadrature set. The contribution to the scattering source can then be computed using a sum over angles of weights times quantities of interest. Typically, iterative schemes from operator splitting are used to treat the scattering (and fission) source terms that arise in this coupled set of partial differential equations [1].
Source iteration (SI), often accompanied by preconditioners or synthetic accelerators, is a common iteration approach: the contribution to the solution from the scattering source lags, while the angular flux is solved in every ordinate direction via “sweeps” through the spatial domain [2]. SI sweeps in Cartesian geometries readily parallelize over the number of angles. While any parallelization improves performance, a scheme that is embarrassingly parallel over the dimension with the greatest number of degrees of freedom—space—would be advantageous, especially on vectorized hardware [3, 4]. In a single spatial dimension, SI is “annoyingly serial” in space and cannot be parallelized.
In higher spatial dimensions, many SN production codes that implement SI use a wavefront marching parallel algorithm known as a Koch–Baker–Alcouff scheme [5], also called “full parallel sweeps.” This algorithm begins a sweep in a spatial location where all cell dependencies are known from boundary information (e.g., a corner). From there, on a hypothetical orthogonal 2D spatial grid the two nearest neighbor cells are computed independently in parallel; the next step would be across four cells. This diagonally expanding wavefront continues to march and can compute quantities of interest in parallel for as many cells that lie on the diagonal sweep step. These sweeps are done on structured or unstructured finite element or finite volume spatial discretization with backward Euler or Crank–Nicholson time stepping iterates. Source iteration is often solved with preconditioned fixed-point (Richardson) or Krylov sub-space methods (e.g., GMRES) [6].
@@
[3] previously studied cell-wise block Jacobi and cell-wise block Gauss–Seidel as a potentially superior iterative scheme over SI preconditioned with diffusion synthetic acceleration on vectorized architectures. They hypothesized that OCI schemes might outperform an SI preconditioned with diffusion synthetic acceleration and using full-parallel sweeps in terms of wall-clock runtime, because of OCI’s parallelism over the dominant domain (space), the ability to take advantage of vendor-supplied LAPACK type libraries, high arithmetic-intensity operations present in an OCI algorithm, and superior spectral properties in the thick limit. Rosa et al. conducted Fourier analysis and implemented OCI in a 2D, multi-group, steady-state code using bilinear discontinuous finite elements to discretize space, a multi-group scheme in energy distribution, and cell-wise block Jacobi and cell-wise block Gauss–Seidel iterations. The study was conducted on the (then) state-of-the-art RoadRunner supercomputer and took advantage of its 64-bit PowerXCell vectorized accelerator. However, the acceleration per iteration in the block Gauss–Seidel OCI implementation did not make up for the degradation of convergence that OCI methods incur in the thin limit.
@@
When solving initial value problems, Crank–Nicholson, backward Euler, or other time-stepping schemes may be employed. When implemented, effective time step size ($\Delta t$) and radiation speed ($v$) are inversely proportional to the effective total cross section ($\Sigma$). Returning to Fig. 1, the total cross section ($\Sigma$) influences both the optical thickness of the cell ($\delta$) and the scattering ratio ($c$), so increasing or decreasing $\Sigma$ will impact convergence behavior. Spectral radius for both iterative methods decreases as the scattering ratio decreases, but the spectral radius of OCI also decreases with increasing optical thickness, which in turn depends on $\Sigma$. When solving optically thin and highly scattering problems, small increases in $\Sigma$ (and for time-dependent problems, decreases in $\Delta t$) may drastically improve the relative performance of OCI compared to SI. This hypothesis motivates our work, along with evaluating cell-wise algorithms on modern GPU accelerators and exploring higher-order space-time discretization schemes.
@@ 2.1 Derivation of space and time discretization for one-cell inversion
Solving Eq. (3) iteratively requires operator splitting. Unknown values (from the current iteration, noted by $l+1$) are moved to the left-hand side to form a large system of linear equations. In SI, the scalar flux in the scattering source is evaluated at the previous iteration $(l)$, decoupling angles and coupling space. OCI allows the fluxes incident to the cell—defined by upstream closures—to lag, thus decoupling cells from one another within an iteration.
@@
$[\mathbf{S}_{j}]_{k.l}=\begin{cases}\frac{\Delta x_{j}\Sigma_{s,j}}{4}w_{|(l-k)|/3)},&\text{if $\mod{\frac{(l-k)}{3}=0}$}\\ 0,&\text{otherwise}\end{cases}\;,$ (10)
where $w$ are the angular quadrature weights. Finally,
@@ 2.2 Fourier analysis: time-stepping scheme
where $\lambda_{K}$ are the eigenvalues of $\bm{K}$ [17, 16]. $\bm{K}$ can be derived for a given model problem. We consider a model problem consisting of a homogeneous infinite medium with no scattering to derive the eigenfunction of the time-dependent multiple balance, simple corner balance discretization scheme. Since this problem has no scattering, each angle can be solved independently of every other angle, and no operator splitting is required. We first start by describing the absolute error of the angular flux at time step $(k+1/2)$:
@@
Then, we define a Fourier ansätz for the error propagated through a time step:
$f_{k+1/2,j,L}$ $=\lambda^{k+1}a_{t}e^{i\omega j}\;,$ $f_{k+1/2,j,R}$ $=\lambda^{k+1}b_{t}e^{i\omega j}\;,$ (17a)
$f_{k,j,L}$ $=\lambda^{k}c_{t}e^{i\omega j}\;,$ $f_{k,j,R}$ $=\lambda^{k}d_{t}e^{i\omega j}\;,$ (17b)
$f_{k-1/2,j,L}$ $=\lambda^{k}a_{t}e^{i\omega j}\;,$ $f_{k-1/2,j,R}$ $=\lambda^{k}b_{t}e^{i\omega j}\;,$ (17c)
where $k$ is the time step, $i=\sqrt{-1}$, $\lambda$ is the eigenvalue, $\omega$ is the wave number, $j$ is cell index and $a_{t}$, $b_{t}$, $c_{t}$, and $d_{t}$ are components of the eigenvector. Substituting our ansätz into the error form of Eqs. (3) and assuming $\mu>0$, we simplify to form:
$\frac{\Delta x}{2}\frac{1}{v\Delta t}(\lambda a_{t}-a_{t})+\mu\left(\frac{c_{t}+d_{t}}{2}-d_{t}e^{-i\omega}\right)+\frac{\Sigma\Delta x}{2}c_{t}=0\;,$ (18a)
$\frac{\Delta x}{2}\frac{1}{v\Delta t}(\lambda b_{t}-b_{t})+\mu\left(c_{t}e^{i\omega}-\frac{c_{t}+d_{t}}{2}\right)+\frac{\Sigma\Delta x}{2}d_{t}=0\;,$ (18b)
$\frac{\Delta x}{2}\frac{2}{v\Delta t}(\lambda a_{t}-c_{t})+\lambda\mu\left(\frac{a_{t}+b_{t}}{2}-b_{t}e^{-i\omega}\right)+\frac{\Sigma\Delta x}{2}a_{t}\lambda=0\;,$ (18c)
$\frac{\Delta x}{2}\frac{2}{v\Delta t}(\lambda b_{t}-d_{t})+\lambda\mu\left(d_{t}e^{i\omega}+\frac{c_{t}+d_{t}}{2}\right)+\frac{\Sigma\Delta x}{2}b_{t}\lambda=0\;,$ (18d)
combining Eq. (18a) into (18b):
$\begin{bmatrix}c_{t}\\ d_{t}\end{bmatrix}=\bm{K_{+}}^{-1}\frac{\Delta x}{2}\frac{1}{v\Delta t}(1-\lambda)\begin{bmatrix}a_{t}\\ b_{t}\end{bmatrix}\;,$ (19)
@@
$\lambda\left(\bm{K_{+}}+\frac{\Delta x}{v\Delta t}\bm{I}\right)\begin{bmatrix}a_{t}\\ b_{t}\end{bmatrix}=\frac{\Delta x}{v\Delta t}\begin{bmatrix}c_{t}\\ d_{t}\end{bmatrix}\;,$ (21)
@@
$\lambda\begin{bmatrix}a_{t}\\ b_{t}\end{bmatrix}=\left[\bm{K_{+}}+\frac{\Delta x}{v\Delta t}\bm{I}+\gamma\bm{K_{+}}^{-1}\right]^{-1}\gamma\bm{K_{+}}^{-1}\begin{bmatrix}a_{t}\\ b_{t}\end{bmatrix}\;,$ (22)
@@
$\bf{a}=\begin{bmatrix}a_{t},&b_{t}\end{bmatrix}^{T}\;.$ (25)
@@
This system can be numerically solved after making discrete selections of $\mu\in[-1,1]$ and $\omega\in(0,2\pi]$ at a point in the perameter space ($\Delta x$, $\Delta t$, $v$, $\Sigma$) with numpy.max(numpy.abs(numpy.linalg.eig(K)) [19].
@@ 2.3 Fourier analysis: OCI iterative scheme
and our Fourier anzats on a functional form of that error
@@
Now, substitute the ansätz and upstream closures into the error form of Eq. (3) and derive the eigensystem. This is done by (1) collecting like terms, (2) dividing both sides by $\omega^{(l)}e^{i\Sigma x_{j}}$, (3) isolating terms with a remaining $\omega$ to the left-hand side, and finally (4) forming the eigensystem into the iteration matrix over all angular directions:
@@ 2.4 OCI multi-group transport
We extend our single energy derivations presented in Section 2.1 to be energy dependent. Elements of the $\mathbf{S_{g^{\prime}\rightarrow g}}$ matrix are now defined by
$[\mathbf{S}_{g^{\prime}\rightarrow g,j}]_{k.l}=\begin{cases}\frac{\Delta x_{j}\Sigma_{s,g^{\prime}\rightarrow g,j}}{4}w_{|(l-k)|/3)},&\text{if $\mod{\frac{(l-k)}{3}=0}$}\\ 0,&\text{otherwise}\end{cases}\;,$ (39)
where $g^{\prime}\rightarrow g$ indicates transfer from group $g^{\prime}$ to group $g$ and $w$ are the quadrature weights. The systems described in Section 2.1 now include multi-group scattering:
$\bm{A}_{j}=\begin{bmatrix}\bm{L}_{c,j,1}-\bm{S}_{1\rightarrow 1,j}&-\bm{S}_{2\rightarrow 1,j}&\cdots&\cdots&-\bm{S}_{G\rightarrow 1,j}\\ -\bm{S}_{1\rightarrow 2,j}&\ddots&&&\vdots\\ \vdots&&\bm{L}_{c,j,g}-\bm{S}_{g\rightarrow g,j}&&\vdots\\ \vdots&&&\ddots&\vdots\\ -\bm{S}_{1\rightarrow G,j}&\cdots&\cdots&&\bm{L}_{c,j,G}-\bm{S}_{G\rightarrow G,j}\end{bmatrix}\;,$ (40)
@@
$\bm{b}_{j}=\begin{bmatrix}b_{j,1},&b_{j,2},&\cdots,&b_{j,G}\end{bmatrix}^{T}\;.$ (41)
@@
in this work we use dense representations in each cell because the matrix memory size for 1D transport is not limiting.

### v2

@@ Abstract
To find deterministic solutions to the transient discrete-ordinates neutron-transport equation, source iterations (SI) are typically used to lag the scattering (and fission) source terms from subsequent iterations. For Cartesian geometries in one dimension, SI is parallel over the number of angles but not spatial cells; this is a disadvantage for many-core compute architectures like graphics processing units. One-cell inversion (OCI) is a class of alternative iterative methods that allow space-parallel radiation transport on heterogeneous compute architectures. For OCI, previous studies have shown that, in steady-state computations, spectral radius tends to unity when cells are optically thin regardless of the scattering ratio. In this work, we analyze how the convergence rate of an OCI scheme behaves when used for time-dependent neutron transport computations. We derive a second-order space-time discretization method from the simple corner balance and multiple balance time discretization schemes and show via Fourier analysis that it is unconditionally stable through time. Then, we derive and numerically solve the Fourier systems for both OCI and SI splittings of our discretization, showing that small mean-free times improve the spectral radius of OCI more than SI, and that spectral radius for OCI tends to zero as mean free time gets smaller. We extend both solvers to be energy dependent (using the multi-group assumption) and implement on an AMD MI250X using vendor-supplied batched LAPACK solvers. Smaller time steps improve the relative performance of OCI over SI, and, even when OCI requires more iterations to converge a problem, those iterations can be done much faster on a GPU. This leads to OCI performing better overall than SI on GPUs.
@@ 1 Introduction
Simulating transient particle transport is often required when computing the solution to a number of multi-physics problems, including burst criticality experiments, fission reactor accidents, and other highly dynamic systems. Finding deterministic solutions to the transient neutron transport equation requires some method of treating the contribution of scattering described by an integral. This is either done by taking moments of the neutron transport equation and making a closure assumption, or by using quadrature to discretize the integral over angle. The latter is called the method of discrete ordinates (or SN method, where $N$ is the number of angles in a one-dimensional quadrature set) that forms a coupled set of simultaneous PDEs, with one for every direction in a given quadrature set. The contribution to the scattering source can then be computed using a sum over angles of weights times quantities of interest. Typically, iterative schemes from operator splitting are used to treat the scattering (and fission) source terms that arise in this coupled set of partial differential equations [1].
Source iteration (SI), often accompanied by preconditioners or synthetic accelerators, is a common iteration approach: the contribution to the solution from the scattering source lags, while the angular flux is solved in every ordinate direction via “sweeps” through the spatial domain [2]. SI sweeps in Cartesian geometries readily parallelize over the number of angles. While any parallelization improves performance, a scheme that is embarrassingly parallel over the dimension with the greatest number of degrees of freedom—space—would be advantageous, especially on vectorized hardware [3, 4]. In slab geometry, SI sweeps can be parallelized in angle and energy groups (via Jacobi iteration), but are serial in space as information about the angular flux incident on edges of each cell is required before the computation can proceed.
In higher spatial dimensions, many SN production codes that implement SI use a wavefront marching parallel algorithm known as a Koch–Baker–Alcouffe scheme [5], also called “full parallel sweeps.” This algorithm begins a sweep in a spatial location where all cell dependencies are known from boundary information (e.g., a corner). From there, on a hypothetical orthogonal 2D spatial grid the two nearest neighbor cells are computed independently in parallel; the next step would be across four cells. This diagonally expanding wavefront continues to march and can compute quantities of interest in parallel for as many cells that lie on the diagonal sweep step. These sweeps are done on structured or unstructured finite element or finite volume spatial discretizations with backward Euler or Crank–Nicholson time stepping iterations. Source iteration is often solved with preconditioned fixed-point (Richardson) or Krylov sub-space methods (e.g., GMRES) [6].
@@
[3] previously studied cell-wise block Jacobi and cell-wise block Gauss–Seidel as a potentially superior iterative scheme over SI preconditioned with diffusion synthetic acceleration on vectorized architectures. They hypothesized that OCI schemes might outperform an SI preconditioned with diffusion synthetic acceleration and using full-parallel sweeps in terms of wall-clock runtime, because of OCI’s parallelism over the dominant domain (space), the ability to take advantage of vendor-supplied LAPACK type libraries, high arithmetic-intensity operations present in an OCI algorithm, and superior spectral properties in the thick limit. Rosa et al. conducted Fourier analysis for and implemented OCI in a 2D, multi-group, steady-state code using bilinear discontinuous finite elements to discretize space. They paired this with either a cell-wise block Jacobi and cell-wise block Gauss–Seidel iteration algorithm. The study was conducted on the (then) state-of-the-art RoadRunner supercomputer and took advantage of its 64-bit PowerXCell vectorized accelerator. However, the acceleration per iteration in the block Gauss–Seidel OCI implementation did not make up for the degradation of convergence that OCI methods incur in the thin limit.
@@
When employing implicit time-differencing schemes (e.g., Crank–Nicholson, backward Euler), each time step involves the solution of a steady-state transport problem with an effective total cross section that includes a time absorption term, proportional to $1/(v\Delta t)$, where $\Delta t$ is time step size and $v$ is radiation speed. Returning to Fig. 1, the macroscopic total cross section ($\Sigma$) influences both the optical thickness of the cell ($\delta$) and the scattering ratio ($c$), so increasing or decreasing $\Sigma$ will impact convergence behavior. Spectral radius for both iterative methods decreases as the scattering ratio decreases, but the spectral radius of OCI also decreases with increasing optical thickness, which in turn depends on $\Sigma$. When solving optically thin and highly scattering problems, small increases to $\Sigma$ (and for time-dependent problems, decreases in $\Delta t$ or $v$) may drastically improve the relative performance of OCI compared to SI. This hypothesis motivates our work, along with evaluating cell-wise algorithms on modern GPU accelerators and exploring higher-order space-time discretization schemes.
@@ 2.1 Derivation of space and time discretization for one-cell inversion
Solving Eqs. (3) iteratively requires operator splitting. Unknown values (from the current iteration, noted by $(l+1)$) are moved to the left-hand side to form a large system of linear equations. In SI, the scalar flux in the scattering source is evaluated at the previous iteration $(l)$, decoupling angles and coupling space. OCI allows the fluxes incident to the cell—defined by upstream closures—to lag, thus decoupling cells from one another within an iteration.
@@
$[\mathbf{S}_{j}]_{k,l}=\begin{cases}\frac{\Delta x_{j}\Sigma_{s,j}}{4}w_{\frac{|(r-s)|}{3}},&\text{if $\mod{\frac{(r-s)}{3}=0}$}\\ 0,&\text{otherwise}\end{cases}\;,$ (10)
where $w$ are the angular quadrature weights, and $r$ and $s$ are the rows and columns of the scattering matrix, respectively. Finally,
@@ 2.2 Fourier analysis: time-stepping scheme
where $\lambda_{K}$ are all the eigenvalues of $\bm{K}$ [17, 16]. $\bm{K}$ can be derived for a given model problem. We consider a model problem consisting of a homogeneous infinite medium with no scattering to derive the eigenfunction of the time-dependent multiple balance, simple corner balance discretization scheme. Since this problem has no scattering, each angle can be solved independently of every other angle, and no operator splitting is required. We first start by describing the absolute error of the angular flux at time step $(k+1/2)$:
@@
We then assume that each unknown representing the error on our discretization stencil is expanded in a series of temporal Fourier modes, where each mode has a coefficient ($a$, $b$, $c$, or $d$), an amplitude ($\lambda$), and a shape function ($e^{i\omega x}$). We can then define a Fourier ansatz for the error propagated through a time step:
$f_{k+1/2,j,L}$ $=\lambda^{k+1}ae^{i\omega j}\;,$ $f_{k+1/2,j,R}$ $=\lambda^{k+1}be^{i\omega j}\;,$ (17a)
$f_{k,j,L}$ $=\lambda^{k}ce^{i\omega j}\;,$ $f_{k,j,R}$ $=\lambda^{k}de^{i\omega j}\;,$ (17b)
$f_{k-1/2,j,L}$ $=\lambda^{k}ae^{i\omega j}\;,$ $f_{k-1/2,j,R}$ $=\lambda^{k}be^{i\omega j}\;,$ (17c)
where $k$ is the time step, $i=\sqrt{-1}$, $\lambda$ is the eigenvalue, $\omega$ is the wave number, and $j$ is cell index. Substituting our ansatz into the error form of Eqs. (3) and assuming $\mu>0$, we simplify to form
$\frac{\Delta x}{2}\frac{1}{v\Delta t}(\lambda a-a)+\mu\left(\frac{c+d}{2}-de^{-i\omega}\right)+\frac{\Sigma\Delta x}{2}c=0\;,$ (18a)
$\frac{\Delta x}{2}\frac{1}{v\Delta t}(\lambda b-b)+\mu\left(ce^{i\omega}-\frac{c+d}{2}\right)+\frac{\Sigma\Delta x}{2}d=0\;,$ (18b)
$\frac{\Delta x}{2}\frac{2}{v\Delta t}(\lambda a-c)+\lambda\mu\left(\frac{a+b}{2}-be^{-i\omega}\right)+\frac{\Sigma\Delta x}{2}a\lambda=0\;,$ (18c)
$\frac{\Delta x}{2}\frac{2}{v\Delta t}(\lambda b-d)+\lambda\mu\left(de^{i\omega}+\frac{c+d}{2}\right)+\frac{\Sigma\Delta x}{2}b\lambda=0\;.$ (18d)
Next, we combine Eq. (18a) into (18b):
$\begin{bmatrix}c\\ d\end{bmatrix}=\bm{K_{+}}^{-1}\frac{\Delta x}{2}\frac{1}{v\Delta t}(1-\lambda)\begin{bmatrix}a\\ b\end{bmatrix}\;,$ (19)
@@
$\lambda\left(\bm{K_{+}}+\frac{\Delta x}{v\Delta t}\bm{I}\right)\begin{bmatrix}a\\ b\end{bmatrix}=\frac{\Delta x}{v\Delta t}\begin{bmatrix}c\\ d\end{bmatrix}\;,$ (21)
@@
$\lambda\begin{bmatrix}a\\ b\end{bmatrix}=\left[\bm{K_{+}}+\frac{\Delta x}{v\Delta t}\bm{I}+\gamma\bm{K_{+}}^{-1}\right]^{-1}\gamma\bm{K_{+}}^{-1}\begin{bmatrix}a\\ b\end{bmatrix}\;,$ (22)
@@
$\bf{a}=\begin{bmatrix}a,&b\end{bmatrix}^{T}\;.$ (25)
@@
This system can be numerically solved after making discrete selections of $\mu\in[-1,1]$ and $\omega\in(0,2\pi]$ at a point in the perameter space ($\Delta x$, $\Delta t$, $v$, $\Sigma$) with numpy.max(numpy.abs(numpy.linalg.
eig(K)))
 [19].
@@ 2.3 Fourier analysis: OCI iterative scheme
and our Fourier ansatz on a functional form of that error
@@
Now, substitute the ansatz and upstream closures into the error form of Eq. (3) and derive the eigensystem. This is done by (1) collecting like terms, (2) dividing both sides by $\omega^{(l)}e^{i\Sigma x_{j}}$, (3) isolating terms with a remaining $\omega$ to the left-hand side, and finally (4) forming the eigensystem into the iteration matrix over all angular directions:
@@ 2.4 OCI multi-group transport
We extend our single-energy derivations presented in Section 2.1 to be energy dependent. Elements of the $\mathbf{S_{g^{\prime}\rightarrow g}}$ matrix are now defined by
$[\mathbf{S}_{g^{\prime}\rightarrow g,j}]_{k.l}=\begin{cases}\frac{\Delta x_{j}\Sigma_{s,g^{\prime}\rightarrow g,j}}{4}w_{\frac{|(r-s)|}{3}},&\text{if $\mod{\frac{(r-s)}{3}=0}$}\\ 0,&\text{otherwise}\end{cases}\;,$ (39)
where $g^{\prime}\rightarrow g$ indicates transfer from group $g^{\prime}$ to group $g$ and $w$ are the quadrature weights.
The full system of linear equations in all groups and angles in cell $j$ becomes
$\bm{A}_{j}\bm{\Psi}_{j}=\bm{b}_{j}\;,$ (40a)
@@
$\bm{A}_{j}=\begin{bmatrix}\bm{L}_{c,j,1}-\bm{S}_{1\rightarrow 1,j}&-\bm{S}_{2\rightarrow 1,j}&\cdots&\cdots&-\bm{S}_{G\rightarrow 1,j}\\ -\bm{S}_{1\rightarrow 2,j}&\ddots&&&\vdots\\ \vdots&&\bm{L}_{c,j,g}-\bm{S}_{g\rightarrow g,j}&&\vdots\\ \vdots&&&\ddots&\vdots\\ -\bm{S}_{1\rightarrow G,j}&\cdots&\cdots&&\bm{L}_{c,j,G}-\bm{S}_{G\rightarrow G,j}\end{bmatrix}\;,$ (40b)
and Eqs. (9) and (11) are extended to multi-group by
$\bm{b}_{j}=\begin{bmatrix}b_{j,1},&b_{j,2},&\cdots,&b_{j,G}\end{bmatrix}^{T}$ (40c)
and
$\bm{\Psi}_{j}=\begin{bmatrix}\Psi_{j,1},&\cdots&\Psi_{j,g},&\cdots,&\Psi_{j,G}\end{bmatrix}^{T}\;,$ (40d)
with otherwise similar structure.
@@
which approaches 25% with large angular and group counts, in this work we use dense representations in each cell because the matrix memory size for 1D transport is not limiting.